OpenStudy (anonymous):

Find the equation of the line perpendicular to the graph of −4x−y=4 that contains the point (−2,−2).

OpenStudy (anonymous):

y+2=-4(x+2)

OpenStudy (gitface):

Begin by rearranging your equation so that it is in the form \[y=mx+b\] From there, determine your slope, namely the value of m. The slope of a perpendicular line is the negative reciprocal of the slope of the original line. Using that, and your point (-2,-2), use the point slope formula \[y-y_{1}=m(x-x _{1})\] Where m is your slope of the perpendicular line and y1 and x1 are your x and y values given in your point on the graph. Solve the given expression for y.

OpenStudy (anonymous):

-4x-4=y slope of dis line is -4.which is m1. for perpencilar lined condtion is, m1*m2=-1 -4*m2=-1 m2=1/4

OpenStudy (anonymous):

dis is the slope of required equation bcoz the lines are perpendicular.if they were parallel,slopes would have benn same.now eqn will be, (y-y1)=m(x-x1) where m is the slope which we found out

OpenStudy (anonymous):

(y-(-2))=1/4(x-(-2)) y+2=1/4(x+2) 4(y+2)=x+2 4y+8=x+2 4y-x=2-8 4y-x=-6

OpenStudy (anonymous):

thanks